OpenStudy (anonymous):

evaluate teh polynomial for x=-1 2x2-3x+8 would the answer be 7

OpenStudy (anonymous):

i mean 2x^2-3x+8

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[2(-1)^2-3(-1)+8=2\cdot1+3+8=13\]

OpenStudy (anonymous):

2(-1)^2 - 3 (-1) + 8 2 - (-3) + 8 = 13

OpenStudy (anonymous):

how you got 7? or it was just question?

OpenStudy (anonymous):

i dont even know i think i 8-1

OpenStudy (anonymous):

you need to put -1 instead of x